OpenStudy (anonymous):

Find the product of (x + 5)^2

OpenStudy (anonymous):

\[(x+y)^2=x^2+2xy+y^2\]

OpenStudy (anonymous):

general formula, can apply to all of these problems you are posting

OpenStudy (anonymous):

\[(x-y)^2=x^2-2xy+y^2\] for subtraction

OpenStudy (anonymous):

I think that's way too abstract for her to understand. :/

OpenStudy (anonymous):

Thanks(: